### v2

# FocusFlow: 3D Gaze-Depth Interaction in Virtual Reality Leveraging Active Visual Depth Manipulation

Abstract.
Gaze interaction presents a promising avenue in Virtual Reality (VR) due to its intuitive and efficient user experience. Yet, the depth control inherent in our visual system remains underutilized in current methods. In this study, we introduce FocusFlow, a hands-free interaction method that capitalizes on human visual depth perception within the 3D scenes of Virtual Reality. We first develop a binocular visual depth detection algorithm to understand eye input characteristics. We then propose a layer-based user interface and introduce the concept of “Virtual Window” that offers an intuitive and robust gaze-depth VR interaction, despite the constraints of visual depth accuracy and precision spatially at further distances. Finally, to help novice users actively manipulate their visual depth, we propose two learning strategies that use different visual cues to help users master visual depth control. Our user studies on 24 participants demonstrate the usability of our proposed virtual window concept as a gaze-depth interaction method. In addition, our findings reveal that the user experience can be enhanced through an effective learning process with adaptive visual cues, helping users to develop muscle memory for this brand-new input mechanism. We conclude the paper by discussing strategies to optimize learning and potential research topics of gaze-depth interaction.
gaze interaction, visual depth, virtual reality, user interface design
journalyear: 2024copyright: acmlicensedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642589isbn: 979-8-4007-0330-0/24/05ccs: Human-centered computing HCI design and evaluation methodsccs: Human-centered computing Virtual realityccs: Human-centered computing Interaction design theory, concepts and paradigms
Figure 1. Concept of FocusFlow. (a) When users hover over an object, a visual cue will appear to guide the users to shift their visual depth between layers at different distances. (b) Users can perform hands-free selection to activate a Virtual Window in front, by pulling their visual depth closer along the z-dimension. (c) Users can exit the Virtual Window by naturally pushing the visual depth further back to the wall.
(a) Hovering Over Target Object
(b) Activation Through Pulling Closer
(c) Exit Through Pushing Further
FocusFlow is a gaze interaction method that enhances traditional methods by incorporating the parallax of the user’s eyes to calculate visual depth. By seamlessly integrating z-dimension information into the interaction logic, users can intuitively shift their gaze point between different layers, enabling hands-free selection.
Figure 1. Concept of FocusFlow. (a) When users hover over an object, a visual cue will appear to guide the users to shift their visual depth between layers at different distances. (b) Users can perform hands-free selection to activate a Virtual Window in front, by pulling their visual depth closer along the z-dimension. (c) Users can exit the Virtual Window by naturally pushing the visual depth further back to the wall.
1. Introduction
Gaze interaction has gained popularity as an input method for 3D interaction in mixed reality (XR) headsets (Sidenmark et al., 2021; Lu et al., 2021a; Adhanom et al., 2023), where the eye movements and gestures would be used as an input. Current gaze-based XR interaction methods often rely on gaze as a means of pointing, supplemented by other inputs such as hand gestures (Yu et al., 2021), head movement (Sidenmark et al., 2021), or dwell time (Penkar et al., 2012; Piumsomboon et al., 2017) for selection confirmation to avoid the so-called ”Midas Touch” problem (Velichkovsky et al., 1997) – the problem of distinguishing intentional gaze inputs from involuntary fixations as the users perceive the scene. However, these works primarily utilize the direction of gaze and overlook the valuable gaze depth information, representing an additional free input dimension along the z-axis (Figure 1). To see how, imagine a window through which you can see scenery outside. By bringing your gaze focus closer to the window, you blur out the distant scenery and are able to perceive dust particles on the window surface instead. Another example is autostereograms (Tyler and Clarke, 1990), where slightly adjusting your visual depth allows you to perceive a 3D image (see example in Appendix A). In both of these examples, users actively manipulate their visual depth to select which object to observe. In this research, we aim to leverage this natural eye behavior and convert it into an intuitive and learnable interaction input in Virtual Reality (VR).
More recent works on gaze-based VR/AR interaction demonstrate the great potential of visual depth as an interaction input to solve the Midas touch problem. These methods either guide the user to look at physical or virtual objects at different depths (Ahn et al., 2020; Pai et al., 2016; Wang et al., 2022a, b; Sidenmark et al., 2023) or rely on voluntary eye convergence and divergence (Kirst and Bulling, 2016; Kudo et al., 2013) by asking the users focusing on the nose or imagining to fixate on some point behind the display plane. However, these works lack an intuitive and systematic User Interface (UI) design to guide the users in manipulating their visual depth, leading to limited application scenarios and potentially user frustration. In addition, most of these works lack a thorough user study to evaluate usability and efficiency of this interaction modality.
In particular, leveraging visual depth as an interaction input is challenging due to several practical factors. First, the natural eye movement is inherently unstable, resulting in noisy visual depth estimates. Moreover, the eye tracking modules in the headsets suffer from limited frame rate and jitters, adding extra noise to depth estimates. So, a usable interaction mechanism is required that is robots against imprecise eye input signal. Second, user’s depth fixation naturally relies on the presence of objects. So, it may feel counter-intuitive if the user is asked to transition to a depth without any visual referent. So, it is crucial to ensure that novice users can intuitively and effortlessly learn and use this interaction and a carefully designed learning procedure is needed to guide the users on how to master this new interaction mechanism. To tackle these challenges, we employed an empirical research approach. Initially, we devised a visual depth detection method for off-the-shelf VR headsets and conducted experiments to investigate the properties of visual depth input. Building upon the insights gained from these experiments, we subsequently developed our proposed user interface and incorporated a learning procedure with in-stages visual cues to help users adapt to the depth-control process.
As such, we present FocusFlow, a novel gaze-depth interaction approach that realizes the concept of “Virtual Window” (just like the examples mentioned above) by leveraging the 3D nature of immersive scenes and the human ability to actively shift between different visual depths. FocusFlow employs the eye trackers in VR headsets to estimate the user’s visual depth. It then defines a layer-based UI design with gaze interaction inputs. Layers are defined as transparent virtual windows at different depths that can be activated through voluntary change of visual depth to where the window is located. More specifically, FocusFlow uses the user’s gaze direction for selecting the object of interest for interaction and user’s visual depth to activate virtual windows that are otherwise invisible in the VR scene (shown in Figure 1). These virtual windows can offer different applications, from a quick preview to activating new control panels, or offering a zoom lens with more information about a selected object.
To master this new interaction mechanism, we incorporate adaptive visual cues in the UI design and a learning process that allows users to establish a clear connection between the eye behavior and visual depth control, facilitating an intuitive and effortless interaction for novice users. The key intuition is that users can develop muscle memory for freely shifting between different focal depths even if empty of objects. We propose two learning strategies, in-stages learning, and adaptive learning. The in-stages learning uses a series of strong to weak visual cues at a certain depth (e.g. where the transparent virtual window is located) so the user can practice visual depth transitions by just moving her gaze from the target object to the cue. While these cues are effective in practicing gaze-depth control, they cause visual interference for the user during non-interaction times. The proposed adaptive learning process addresses this issue by adaptively changing the transparency of visual cues as the user gets closer to the intended interaction depth. The transparency level of the cue also serves as a feedback for the user to better control their focal depth changes. To help master this interaction, FocusFlow gradually reduces the guided depth range until reaching no visual cue.
To assess the usability and learnability of our proposed interaction, we designed both quantitative and qualitative evaluations that take into account factors such as input efficiency, eye behavior, and cognitive load. The evaluation results from 24 participants prove that FocusFlow achieves high input efficiency and intuitive user experience on the hands-free selection tasks. Our design solves Midas Touch problem while retaining a natural interaction process. In addition, the result shows users can easily learn and master the gaze-depth method through our adaptive learning strategy. It is evident that the adaptive visual cues incorporated within our design substantially augmented users’ depth perception during the visual depth transition phase.
To the best of our knowledge, this work is the first to offer an intuitive and efficient gaze-depth interaction method with extensive evaluation. In summary, this presented work makes the following contributions:
  1. (1)
    We extensively analyze the characteristics of gaze-depth estimates and propose strategies for effectively utilizing this imprecise input information.
  2. (2)
    We develop a layer-based user interface for gaze-depth interaction by realizing the concept of a “virtual window”. This interface enables intuitive and robust hands-free VR interaction inputs while avoiding Midas Touch problem.
  3. (3)
    We presented a series of visual cues that provide users with depth perception to intuitively adjust their visual depth for interaction.
  4. (4)
    We present and compare two learning strategies, in-stages learning and adaptive learning, to guide novice users in adapting to depth control and gaze-depth interaction in virtual scenes. This learning process enables users with an effortless interaction by helping them develop muscle memory for this new interaction.
  5. (5)
    We conduct an user study and a case study to assess both the system’s performance and the user’s learning process. The results demonstrate the usability and learnability of our proposed system.
2. Related Work
In this section, we first summarize different eye behaviors and the corresponding input techniques which can be used for gaze interactions. Based on this, we then review the evolution of gaze-based interaction in VR and some representative methods. Finally, we discussed the existing gaze-depth interaction techniques to discover the research gap.
2.1. Eye Behaviors and Input Techniques
Eyelid Functions
Eyelid movements, ubiquitous in daily life, also hold value as input mechanisms. Blinking and winking, rapid opening and closing of the eyelid, serve to distribute tears across the eye’s surface. Despite their inherent simplicity, these actions can be random and noisy, complicating the balance between efficiency and accuracy. Consequently, most input techniques leveraging blinking and winking are geared towards accessibility applications (Królak and Strumiłło, 2012). Another area of interest is the detection of eyelid openness (Arvacheh and Tizhoosh, 2006), a measurable value that can be used to control objects (Yi et al., 2022).
Eye Movements
Eye movements have the potential to serve as input signals, as they can reflect attention direction. Notably, fixation is the act of maintaining a steady gaze on a specific location, while saccade involves quick, darting eye movements that suddenly shift the fixation point. Feit et al. (Feit et al., 2017) explored the utilization of saccades and fixations for accurate interaction input, laying the groundwork for many gaze-based interactions. Smooth pursuit movements allow our eyes to steadily track a moving object. Inspired by this, researchers have proposed several pursuit-based gaze input methods for tasks such as selection (Esteves et al., 2015) and text entry (Zeng and Roetting, 2018; Abdrabou et al., 2019).
Vergence
Vergence is a specific type of eye movement that represents the coordinated movement of both eyes in opposite directions to sustain binocular vision As a result of such eye movement, the user’s visual depth will change. Previous works such as Verge-it (Ahn et al., 2020), employ vergence eye movement as an interaction input to select objects. In these techniques, the user is either asked to follow the motion of an object which results in vergence changes, or to look at their nose to create voluntary vergence control. We build our proposed interaction method upon these eye movements and vergence control systems, but we propose an intuitive user interface design that allows the user to perceive visual depth and actively use it for interaction with minimal cognitive load or visual interferences.
$2.$ Gaze-Based VR Interaction
Gaze-ray-based Interaction. Eye movements were first utilized for VR pointing tasks by Tanriverdi and Jacob (Tanriverdi and Jacob, 2000), resulting in faster interaction compared to hand-based gesture methods. Building on this, Sidorakis et al. (Sidorakis et al., 2015) introduced a VR keyboard input method that employs gaze pointing, leading to fewer typing errors than conventional interfaces. Mardanbegi et al. extended this approach with EyeSeeThrough (Mardanbegi et al., 2019), a VR selection method that aligns the intended object with a tool menu using the user’s gaze ray. Piumsomboon et al. (Piumsomboon et al., 2017) further explored the application of gaze direction in VR, proposing different gaze ray selection methods that only leverage the direction of gaze. However, a direction-only gaze selection method causes a so-called Midas Touch problem (Jacob, 1990), where unintentional inputs might be recognized as commands.
Dwell-time-based Gaze Interaction. Researchers have attempted to mitigate Midas Touch problem by incorporating dwell time (Lu et al., 2021c; Lu and Bowman, 2021; Majaranta et al., 2009; Mott et al., 2017; Mardanbegi et al., 2019; Pfeuffer et al., 2020) or similar attention patterns (Piumsomboon et al., 2017). While these dwell-based gaze selection techniques can reduce false triggers, they also slow down the input response due to the added time required for dwell validation, which degrades the interaction efficiency and user experience.
Multi-modal Interaction. Another approach for addressing the Midas Touch problem is to expand the input dimension such as leveraging multi-modal inputs, which combines gaze with another input modality. The typical multi-modal inputs combine the gaze direction and controller inputs (Sendhilnathan et al., 2022; Mardanbegi et al., 2019). Other works combine gaze and head information. For example, Sidenmark et al. proposed Radi-Eye (Sidenmark et al., 2021), which combines head and eye movement to validate selections. Wei et al. (Wei et al., 2023) proposed a prediction method for target selection in AR based on the distribution of eye and head endpoints. More recent papers explore the combination of gaze and pinch (Pfeuffer et al., 2017), gaze and hand alignment(Lystbæk et al., 2022), humming (Hedeshy et al., 2021), or other hand gestures (Yu et al., 2021; Lu et al., 2021a, b; Hou et al., 2023; Bao et al., 2023). In these systems, the pointing information is extracted from eye movement and selection or confirmation of information are done with hand gesture. The gaze and hand multi-modal interactions are also built into commercial products such as Apple Vision Pro. However, involving other modalities inevitably increases the interaction complexity and may not be applicable in certain applications that require hands-free interaction.
An alternative way to expand interaction dimension is to leverage more eye-input information. For instance, unconventional blinking (Lu et al., 2021a; Rebsdorf et al., 2023; Yu et al., 2021) is used as a complementary gaze input to direction, but the proposed technique suffers from unintentional triggering. Yi et al. introduced DEEP (Yi et al., 2022) which leverages eyelid movement as a new input for controlling to allow users to adjust the visual depth in a scene, thus improving the gaze-pointing accuracy when some objects are occluded. However, this approach relies on dwell time, which still suffers from efficiency and usability problems. In this study, we incorporate visual depth as a novel input dimension in gaze interaction to address the Midas Touch problem in gaze selection, while maintaining interaction efficiency.
Gaze-depth-based Interaction. Gaze vergence is an emerging eye-input information that has been used for VR interaction. Compared to multi-modal gaze interaction methods that leverage handheld controllers or hand gestures, visual depth offers a completely hands-free interaction mechanism. Previous works have shown the feasibility of visual depth estimations in the VR headsets by using binocular disparity and stereoscopic vergence in VR headset (Hirzle et al., 2019; Vidal et al., 2014). The concept of visual depth as a new interaction input has explored in the previous works, either by defining a semi-transparent window at a different focal depth (Pai et al., 2016; Wang et al., 2022a, b), tracking voluntary vergence movements (Kirst and Bulling, 2016; Ahn et al., 2020), or matching the vergence changes with the depth changes of a moving objects (Sidenmark et al., 2023) in VR. However, all of these methods lack an end-to-end UI design to guide the users how to actively manipulate their visual depth. As a result, their users suffered from fatigue and difficulty to master this new interaction. In addition, they lack sufficient user studies to evaluate the usability and efficiency of this interaction mechanism in the context of different applications. This paper targets these limitations and proposes a layer-based UI design to leverage active visual depth manipulation despite unstable depth estimates of existing eye tracking systems. In addition, we propose an integrated learning process that leverages visual cues to help users learn how to voluntarily shift their focal depth. We also perform multiple user studies to evaluate user experience and mental model in our gaze-depth interaction process.
2.3. Immersive Gaze-based UI Design
There is a flurry of work on user interface design that incorporate gaze input in VR and AR systems. For example, Pfeuffer et al. (Pfeuffer et al., 2021; Blattgerste et al., 2018) proposes a design space for gaze-input based interfaces for AR applications. Hirzle et al. (Hirzle et al., 2019) introduced a design space for head mounted devices focusing on human depth perception and technical issues. In this paper, we propose a layer-based UI design for virtual reality that integrates active visual depth manipulation and gaze movement for VR interaction. We fully evaluate this new interaction and demonstrate the potential of this gaze-only interaction method as an effortless and intuitive method for everyday VR through a learning process.
3. Examining Gaze Depth Characteristics
In order to design an intuitive and robust gaze-depth interaction, we first conducted a user study to explore the performance of built-in eye trackers in commodity VR headsets and visual depth characteristics.
(a) Visual depth calculation. Two gaze rays are projected onto the x-z plane. The visual depth is calculated as the distance between the intersection of projected gaze rays and the midpoint of the user’s eyes.
(b) Gaze-depth Sensitivity analysis. For a fixed interpupillary distance, a small change in gaze direction at larger angles (aka larger depths) create much larger changes in visual depth. Therefore, larger visual depths are more prone to noises.
Figure 2. Depth calculation and sensitivity analysis.
3.1. Visual Depth Estimation
In order to calculate the user’s visual depth, we use the built-in eye trackers in VR headsets, which capture the positions of eyeballs and converts them into binocular gaze rays originating from eyes and intersecting at a specific depth (see Figure 1(a)). However, two rays in 3 dimensions may not intersect at a point specially with a small error in gaze direction estimations. To address this issue, we leverage the projection of these two rays onto a plane which ensures the intersection of the two rays. We define the projection plane along the two gaze ray origins (x-axes) and z-axes. As such, the distance between the eye line and the intersection point in this projection plane will be equal to the visual depth. To deal with random eye movements (Mughrabi et al., 2022; Zhang, 2021) and the resulting depth jitters, we apply a moving average technique with 0.2 seconds of time window to filter out the high frequency depth fluctuations in the depth estimates.
Figure 3. Experiment settings. The participants are required to look at three static objects located at 0.5m, 1.0m, and 2.0m away from themselves. Then they are asked to follow a moving target moving back and forth between 0.5m and 2.5m away with their gaze.
3.2. Experiment Design For Pilot Study
[…]
We designed two sub-tasks in this experiment. In the first task, participants are asked to stare at a static VR target for five seconds located straight ahead at different distances from the user (0.5m, 1.0m, and 2.0m respectively). The target will be highlighted to capture participants’ attention. In the second task, they are asked to look at a moving target that moves back and forth toward the user at depths between 0.5m to 2.5m. The participants are allowed to blink and take a rest if they feel tired in both stages. Figure 3 shows all the tasks in the experiment.
Figure 4. Visual depth detection results. (a) per-participant histogram of estimated depths for three static targets at different distances; (b) snapshot of estimated depth fluctuations for the moving target with and without de-noising.
$3.$ Takeaways
Figure 4 summarizes the results of the visual depth estimates and visual depth behaviors for the 7 participants. Figure 4a shows the histogram of estimated depths for task 1 with static targets at different depth, represented by different colors. Figure 4b shows snapshots of depth estimates in task 2 (moving target) with and without denoising. We conclude the following visual depth characteristics:
  1. C.1
    Visual depth is a noisy but usable signal: In Figure 4a, we can see that estimated visual depths are always presented as a distribution around the target’s groundtruth depth. This verifies the unavoidable random eye movements even when looking at a single static object. Nonetheless, we are able to find a relatively clear demarcation between the blue distribution (target at 0.5m distance) and the red distribution (target at 2m distance). This suggests that visual depth can be distinguished at relatively coarse granularity, but it is still a usable signal.
  2. C.2
    Estimated Visual depth accuracy decreases with the increase of distance: By comparing the visual depth patterns in Figure 4a and b at different distances, we can conclude that visual depth can be estimated with high accuracy (absolute depth values) and precision (standard deviation of depth distributions) in short distances, but both the depth estimate variations and accuracy drops for larger depths. We then conduct a theoretical sensitivity analysis of the visual depth to the eye rotation and gaze direction. As shown in Figure 1(b), the visual depth becomes more sensitive to the eye rotation changes at larger eye rotation values. So, at longer distances (more tahn 2 meters away), even a small amount of noise or error in gaze direction estimates can result in much larger error in visual depth calculations.
  3. C.3
    De-noising method plays a critical role. The raw depth estimation curve (red) in Figure 4b shows persistent fluctuations, especially when the target distance is larger than one meter from the user. After applying our de-noising method, the new depth curve (red) is smoother and drastic depth fluctuations due to random eye movements are eliminated. This indicates that our de-noising method plays a crucial role in pre-processing the input visual depth data.
3.4. Design Rationales Based on Gaze Depth Characteristics
Our pilot study for characterizing the visual depth estimation sets certain trade-offs on leveraging gaze depth as an interaction input. In one hand, gaze depth estimates are mainly reliable in short distances, with larger error in estimated depths as the distance increases. So, for long distance operations, we can only monitor obvious depth shifts which limites the gaze-depth inputs to coarse grained depth changes only. On the other hand, gaze-based interaction is more useful in large virtual environments for interacting with distant objects, where other modalities (e.g. hand or head gestures) cause user fatigue (Yu et al., 2021). So, the main objective of this work is to break this trade-off by taking into account users’ spatial perception, input efficiency, and user’s learning ability. As such, we define three design goals for FocusFlow. Using FocusFlow, :
  1. D.1
    Users should be able to achieve hands-free selection efficiently and accurately, without encountering the Midas Touch problem. The efficiency of a gaze-based interaction is measured by the wait time between the user input and the system response and is expected to be below the user’s perceptible limit. The accuracy is indicated by the success rate and false trigger rate of an operation (e.g. selection), which means the selection and deselection intention should be accurately captured by the system.
  2. D.2
    Users should be able to perceive the depth and accordingly adjust their focal depth for interaction. Despite the difficulty of users in adjusting their visual depth to a certain length without any reference in space, perceiving the depth by using alternative visual cues can improve voluntary gaze depth changes. The system should offer an intuitive mechanism to grasp the concept of visual depth transitions.
  3. D.3
    Users should be able to develop muscle memory to provide efficient and effortless interaction. An immersive virtual scene should be compatible with user instincts to avoid user discomfort (Weech et al., 2019). Since voluntary gaze-depth change is a brand-new input method for general VR users, learnability is critical to our interaction approach. Some guidance and feedback should be provided to facilitate the user’s learning process.
4. FocusFlow: System Overview
This paper proposes FocusFlow, which consists of three main components to satisfy the design goals mentioned in the previous section:
  • Layer-based UI Design: FocusFlow presents a layer-based UI design for virtual reality applications that supports gaze-depth interaction. In this design, interactive virtual contents are organized into distinct spatial layers at different visual depths relative to the user. The visibility indexes of these layers are defined based on the user’s gaze depth, transitioning between completely opaque when the user’s gaze depth matches with the depth of a layer to completely transparent otherwise. Section 5 explains how FocusFlow avoids Midas Touch problem by carefully selecting the layers in the UI and active gaze-depth manipulation to activate a layer.
  • Gaze-depth Interaction Design: FocusFlow leverages the proposed layer-based UI design and offers a gaze-only interaction mechanism by combining gaze direction and gaze depth. The gaze direction is used for selection of the target object for interaction and gaze depth is used for activating a certain layer in the UI. Section LABEL: discusses a variety of applications that can benefit from this hands-free gaze-only interaction design.
  • Adaptive Interaction Learning Process: FocusFlow takes advantages of a series of visual cues to help users master this brand-new interaction. The visual cues are designed such that they provide users with depth perception and learning opportunities to practice, receive feedback, and master intuitive adjustment of their gaze depth for interaction. Section LABEL: describes two different learning strategies based on in-stages and adaptive visual cues.
5. FocusFlow: Layer-based UI Design
Figure 5. Layer-based UI. The layers are arranged at different depths along the z-direction. The user can match the corresponding layers by changing the visual depth. The matched layers will be activated.
Our vision perception undergoes a significant transformation when there is a change in visual depth. More specifically, a change in the perceived depth causes objects within the original field of view to appear out-of-focus and blurred since they are no longer aligned with our current focal plane. Our goal is to mimic the same intrinsic user experience in the virtual environments by aligning the in-focus objects with the user’s perceived depth. To achieve this, we propose a “layer-based” spatial user interface that organizes virtual content into distinct layers. We also propose an activation logic that defines the visibility index of each layer based on the user’s visual depth. These visibility indexes are dynamically adjusted based on shifts in the user’s visual depth. To create an immersive experience, the layer that is matched with user’s visual depth is visible with low transparency so that the user can perceive its position relative to further layers, while the closer layers are completely invisible. As such, we can enhance the user’s sense of depth perception. Figure 5 shows a high-level overview of the proposed layer-based UI.
Figure 6. Activation logic. (a) Pointing: When the user is looking at the target object in VR environment, the gaze point falls on the object and the visual depth exceeds the activation zone. (b) Depth Shift: When the user’s visual depth falls in the activation zone, the Virtual Window will be activated. (c) The gaze depth change in multiple activations.
5.1. Virtual Window: A Depth-based Selection Widget
Leveraging the proposed layer-based UI, we design Virtual Window, a hands-free selection widget with visual depth input. The high level idea is to replicate the user experience of seeing through a “window” or looking at the “window” itself in the virtual environment. Our vision system is capable of ignoring the presence of window and only focus on scenery behind it. At the same time, by bringing your visual focus closer to the window, you blur out the distant scenery and zoom into dust particles on the window surface instead.
Similarly, this virtual Window is an information layer, floating in front of the user, and it is invisible by default. The user can activate the window by bringing their gaze point closer. When the visual depth is matched with the window distance, the information on the Virtual Window will be activated, as shown in Figure 6. This transition process is highly intuitive, as it provides users a sense of “grabbing in” or “taking a closer look” at detailed information when they bring the visual depth closer. Similarly, to send the window to its default transparent status, the user only need to push their visual depth further, back to the portal layer. The window’s visibility change is a response indicating the activation is successful, and also naturally helps users to maintain the visual depth as they are looking at information on the window.
One challenge with this layer-based UI design is potential false triggers or no triggers due to gaze-depth estimation errors. To avoid false triggers, we employ our findings in the pilot study, described in Section 3, that showed a higher depth estimation accuracy at shorter distances. Therefore, we select the virtual window depth within the 2m range of the user at a distance that no virtual object is present. To avoid false negatives and no triggers, we define an activation zone, shown in Figure 6, which is a predefined spatial range around the actual virtual window depth. If the user’s visual depth falls inside this activation zone, the virtual window will be activated.
Figure 7. Three different visual cues. (a) Strong visual cues appear at the center of the view with constant low transparency. (b) Weak visual cues appear at the margin of the view with constant low transparency. (c) Adaptive visual cues will appear at the center of the view with variable transparency which depends on the user’s gaze depth.
5.2. Visual Cues
While visual depth transition comes naturally to our vision system, it highly relies on the presence of physical/virtual content at that particular depth. So, it may be at first counter-intuitive for the users to perceive the depth and switch their gaze to a point in the space with no visual feature. Previous works on gaze-depth interaction address this issue by asking users to look at their nose to create vergence (Kirst and Bulling, 2016; Ahn et al., 2020). However, this may results in user discomfort and fatigue. Instead, FocusFlow defines three types of visual cues that offer intuitive interaction with minimal visual interference.
The first design is a strong guidance (Figure 7a), which is a static green circle in the center of the user’s view at the close depth. With this, users can transition their visual depth from far to nearby by directly looking at the circle. However, since the circle is always in the user’s field of view, it may be a strong visual interference for some users. The second design is a frame layout in the corners of the view, called weak guidance (Figure 7b). Users can perceive the layer depth through their peripheral vision to transition their visual depth. As such, the strong guidance is a direct referent for the gaze point to focus, but the weak guidance is an indirect referent to the layer depth. The third design is an adaptive guidance (Figure 7c), which is also a circle in the center, but its visibility actively changes based on the user’s visual depth. We map its opacity to a certain range of the user’s visual depth. When users are looking at distant objects, the circle is not visible so there is no visual interference. As users look closer, its opacity increases so users can find the circle and manipulate their gaze depth easier. In this adaptive process, the visual cue is not only a depth referent but also an indicator of the user’s visual depth, which could serve as a feedback to the user; from the opacity or transparency of the circle: the closer the visual depth, the more obvious the circle will be. This helps users understand the gaze-depth interaction mechanism and better perceive the depth information in virtual scenes.
Figure 8. System pipeline of FocusFlow. Eye Tracker provides gaze data to the Interactive Object Selection and Layer Selection modules, and these two modules then send signals to control the activation of the Virtual Window.
6. FocusFlow: Interaction Design
Figure 8 illustrates the overall control logic of FocusFlow’s interaction design. The process is initiated with the eye tracker module in the headsets, which continuously monitors the user’s binocular gaze directions. From the gaze direction information, it can then compute the gaze depth to determine the exact point the user is looking at (as described in Section 3.1). The Eye Tracker also employs a ’moving average’ de-noising technique to refine the gaze data, ensuring that the data is smooth and free from random fluctuations that could lead to depth estimation errors.
The gaze data is then passed to the Interactive Object Selection module, which determines whether the user’s gaze is aligned with an interactive object in the environment. The module mainly searches for intersection of the gaze rays with individual objects in the virtual scene and moves forward to layer selection module if the an object is selected.
Upon selection of the interactive object, FocusFlow continuesly tracks the gaze depth estimates and evaluates whether the visual depth coincides with the activation zone, which is a predefined spatial area where user interaction is intended to occur. If the visual depth falls outside this zone, it is interpreted as a non-interactive state. However, if the gaze depth is within the zone, indicated by a ”Yes” decision, it triggers the interaction signal.
Upon satisfying both the direction and depth conditions, the interaction signal is then conveyed to the Virtual Window Activation module, which is responsible for managing the active window elements on the user interface. It reacts to the interaction signal by activating or deactivating window elements. For instance, if the user’s gaze is directed at and focused on an interactive object within the activation zone, the corresponding virtual window element becomes active with full visibility. Conversely, if the user’s gaze shifts away, the virtual window will disappear and return to an ’Inactive’ state, as depicted in Figure 8.
(a) Quick Preview
(b) Safe Activation
(c) Zoom Lens
Figure 9. Example applications of Virtual Window.
6.1. Interaction Applications
The proposed layer-based UI and interaction offers a flexible design that can support different functionalities for the Virtual Windows. We implement and demonstrate FocusFlow functionality across several example applications:
Quick Preview. Given the expected short depth transition delays in our visual systems (i.e.160-200 milliseconds (Leigh and Zee, 2015)), users can activate the Virtual Window very fast, making it suitable for quick information preview. As Figure 8(a) shows, we can display preview information about the selected item on the virtual window, which is very common in scenarios like gaming, training, 3D creation, virtual tours, etc.
Safe Activation. Compared to actions such as gaze direction change, blinking, and finger movement, wide range visual depth shift is a relatively low-frequency and low-randomness action. Therefore, we can apply the depth activation method to scenarios where we don’t want false touches to occur, which we call “safe activation”. For example, when browsing a web page or watching a video in full screen, we can activate a toolbar by shifting the visual depth. Since our visual depth is relatively stable, we can avoid false touches that could interfere with our viewing experience. (Figure 8(b))
Zoom Lens. Visual depth is an input dimension along z-axis, where occlusion happens. This constraint actually offers logic connection between different layers, for example, the Virtual Window can also be used as zoom lens to magnify the object behind it. (Figure 8(c)) This design connects the information layer (Virtual Window) and portal layer (environment), i.e. providing a zoomed view.
Figure 10. Learning Strategies. In-Stages learning begins with strong guidance at the center of the view, then transfers to weak guidance at the edge, and finally moves to no guidance. Adaptive learning starts with adaptive guidance at the center of the view with variable transparency. As the user progresses, the adaptive range gradually decreases, eventually leading to no guidance.
7. FocusFlow: Interaction Learning Process
Section 5.2 defines several visual cues that could help the users to intuitively shift their gaze depth. However, we argue that depth change is a learnable input method and can become a muscle memory with a proper learning strategy without the need for strong or even any visual cues. To help users develop this new muscle memory, we propose two learning processes based on different series of visual cues: (1) in-stages learning (Figure 10-I), which walks the user through a sequence of strong to weak visual cues, gradually weakening the cue to remove the user reliance on visual cues for gaze depth transitions. (2) adaptive learning (Figure 10-II), which leverages the adaptive visual cue with different transparency indexes for the circular visual cue based on the user’s visual depth. We describe the technical details of these two learning strategies in the next sections.
In-stages Learning During in-stages learning process, the UI walks the user through a sequence of gaze-depth interactions with strong, then weak, and finally no visual cues. The learning process mainly relies on repetition and gradually weakening the cues in a plane dimension. In addition a fixed learning sequence is expected to be used for different users. This learning process is simple and efficient but lacks feedback to users on their depth control and perception quality.
Adaptive Learning The second learning strategy uses the adaptive guidance. At first, a depth range will be defined for activating the visual cue with different transparency indexes assigned to the visual cue based on the user’s visual depth. It should be noted that the visual cue’s depth range is much larger than the activation zone for virtual window activation. While the adaptive transparency of the visual cue serves as a feedback to the user, we further expand this feedback-based learning process by gradually reducing the depth range of the visual cue until we reach the depth at which the Virtual Window is located. In addition to feedback offering, this learning process can be flexibly modified for different users by changing the rate of visual cue’s depth range.
8. User Study
In this section, we focus on two questions: First, can a gaze-depth input method be intuitively learned to the extent of developing a muscle memory? Second, how efficient is our method and what are the advantages compared with the non-depth hands-free methods? To prove the learnability and efficiency of FocusFlow, we conducted a user study with 24 participants, analyzed the effects of two learning strategies, and compared the performance with the baseline dwell-based selection method as detailed below.
8.1. Participants and Apparatus
We recruited two groups of participants from the campus with 12 people for each group (Group 1 and Group 2). All of these 24 individuals only have occasional or no VR experience. Out of these 24 individuals, 6 of them reported slight shortsightedness (3 in Group 1 and 3 in Group 2), while 18 of them had a standard uncorrected vision. All the participants did not wear glasses during the test. In addition, we invited 2 expert users (E1, E2) who had long-term participation in this project as a reference result. The expert users finished all the studies for Group 1 and Group 2. We used the same apparatus as in study 3.2.
(a) Strong Guidance.
(b) Weak Guidance.
(c) Adaptive Guidance.
(d) Activation.
Figure 11. User study in a gallery scenario.
Figure 12. Learning attempts of Group 1. Group 1 adopts the in-stages learning strategy with 5 strong-guidance attempts, 10 weak-guidance attempts and 3 no-guidance attempts.
Figure 13. Learning attempts of Group 2. Group 2 adopts the adaptive learning strategy with 15 adaptive-guidance attempts and 3 no-guidance attempts. Within the 15 adaptive-guidance attempts, the adaptive range will shrink for every 3 attempts, starting with 5 meters and ending with 1 meter. The target object for selection is 8 meters away from the user.
8.2. Procedure
Participants were seated during the experiment, and they could take a rest at any time during the entire process. At first, they put on the headset and performed an eye tracker calibration program. Then they were asked to learn to do gaze selection tasks in a VR gallery shown in Figure 11, which was to select the target painting to the Virtual Window using the gaze-depth interaction method. This learning process is analyzed in Section 8.3.
Then they put on the headset again to test their learning effectiveness and proficiency in operation. Participants were first required to finish 10 selection operations within 5 seconds in the same scene, with the help of visual cues. Then they were required to finish another 10 selection operations within 5 seconds without any visual cues. Once they finished the testing stage, the participants will remove the headset and answer the NASA TLX Workload questionnaires to measure their cognitive load.
Lastly, all the participants in both groups were asked to test the baseline method. They will complete two rounds of 10 selection tasks using the dwell-based gaze interaction method, with time thresholds of 0.5 second and 1 second. Finally, all the participants removed the headset and answered a 7-point Likert scale questionnaire about their user experience. The whole experiment took 20-30 minutes to complete.
8.3. Learning Strategy Analysis
In this section, we discuss the detailed performance of the two proposed two learning procedures and their learning capabilities by comparing the activation time, users’ gaze depth behavior, and the workload during the learning process. With better learning performance and user experience, the adaptive learning strategy is shown as a more effective way for learning gaze-depth interaction than the in-stages learning strategy.
8.3.1. Learning Procedure
Participants in Group 1 are asked to go through the in-stages learning process. As Figure 12 shows, they first do 5 selections with strong guidance, followed by 10 selection attempts with weak guidance, and finally 3 attempts without any visual cues. On the other hand, participants in Group 2 are asked to go through the adaptive learning process. As shown in Figure 13, they perform 15 selection attempts using the adaptive visual cue. The depth range of adaptive guidance start from 5 meters, shrinks 1 meter every 3 attempts, and ends at 1 meter visual depth range which overlaps with the activation zone of Virtual Window. After this adaptive process, the users are asked to perform selections without any visual cue.
(a) Group 1.
(b) Group 2.
Figure 14. Activation time statistics of all the learning attempts for two groups. the central line in the box indicates the median activation time, while the edges of the box show the interquartile range (IQR), which represents the middle 50% of the data. Group 1’s activation times show a wider range with more variability and outliers, suggesting less consistent performance across attempts. Group 2 displays a more consistent activation time across attempts with fewer outliers, indicating a more uniform performance.
8.3.2. Activation Time
We visualize the activation time of all the participant attempts from both groups in the learning stage in Figure 12 and 13, where the color intensity encodes the activation time. Both figures show a trend of decreasing activation time as learning progresses, demonstrated in lighter colors from left to right for each line/participant. We can also see that 7 participants from Group 1 cannot successfully activate the Virtual Window under no guidance setting, while all the participants from Group 2 can successfully do this. This indicates that learning strategy 2 is more capable of helping users form muscle memory during the learning process.
We further analyze the effectiveness of learning by comparing the activation time changes in a statistical way. The box plot in Figure 14 shows the distribution of activation times for each round of attempt within two groups. We can see that the activation time of Group 1 (Figure 13(a)) shows a decreasing trend in median activation time specifically under the strong and weak guidance setting, but this trend does not extend to the no guidance setting. In comparison, Group 2 (Figure 13(b)) shows a consistent decreasing trend under the adaptive guidance and no guidance setting, proving its superiority over learning strategy 1.
8.3.3. Gaze Depth Behavior
We recorded the gaze depth data during the learning process to investigate the effect of visual cues on the interaction behavior. Figure 15 shows the raw gaze depth values of two representative subjects (P8 and P26).
Figure 15. Visual Depth Change. The gaze depth change is visualized as the blue line, and the background color indicates different visual cues settings.
P8 adopted the in-stages learning strategy and experienced strong guidance, weak guidance, and no guidance activation successively. As Figure 15-A shows, the gaze depth transition of P8 is smooth under the strong guidance and weak guidance setting. However, when entering the no guidance setting from weak guidance (Figure 15-B), P8 made several attempts but can not effectively bring his gaze depth close enough to reach the activation threshold. The reason is it’s hard to adjust the gaze depth accurately without any depth referent or muscle memory. The in-stages learning strategy fails to help P8 to develop the muscle memory, so as the system switched to the no guidance setting and there was nothing to be a referent, P8 can not shift his gaze depth to the target distance.
The adaptive learning strategy shows a better effectiveness on the transition from adaptive guidance to no guidance. As Figure 15-C shows, P26 can still stably shift the gaze depth between close distance and far distance, proving the feasibility of forming a muscle memory. This is because the shrinking assisted range is continuously reduce users’ dependence on visual cues, and the changing opacity provides a clear feedback to users to help them understand their input status.
(a) Group 1.
(b) Group 2.
Figure 16. NASA-TLX Results. We measure the subjective workload during two learning process (in-stages learning and adaptive learning) using the NASA-TLX Index. For each entry, lower score means lower workload and better experience.
8.3.4. Workload
We employed the NASA Task Load Index (NASA-TLX) to assess the cognitive and physical workload of two learning processes. The result in Figure 16 shows that the no guidance method imposes higher cognitive and physical demands on the users than the guided methods, indicating the muscle memory development requires some effort. In addition, the cognitive load scores of adaptive guidance is slightly higher than the weak guidance method, because its changing capacity brings more information to users to process. However, it also shows a lower performance load score (lower is better), which means eventually the adaptive guidance can offer better activation performance.
(a) Test with adaptive guidance of Group 2.
(b) Test with no guidance of Group 2.
Figure 17. Activation time for the tests of Group 2 in two guidance settings. In the test with no guidance, users in Group 2 can still finish most of the selection tasks in a similar activation time to the time used in the test with adaptive guidance.
8.4. Performance Analysis
The analysis in Section 8.3 revealed that the adaptive learning strategy enhances participants’ proficiency with the gaze-depth interaction method in FocusFlow. Building on this insight, this section delves into the performance of Group 2 during their test stages, comparing it to the baseline dwell method to assess the efficiency and accuracy of FocusFlow.
8.4.1. Activation Time
We define ’activation time’ as the duration from the moment a user’s gaze targets an object to when the virtual window appears. Figure 16(a) and 16(b) illustrate the activation times across all selection attempts by Group 2 during two tests in the testing stage. It’s important to note that dark squares marked as ’N/A’ represent instances where participants didn’t complete the selection task within the 5-second limit. Remarkably, participants in Group 2 successfully completed the selection tasks in the no-guidance test setting. This achievement underscores their high proficiency with the gaze-depth interaction method, particularly when contrasted with the last several attempts shown in Figure 12 for the in-stage learning strategy. Additionally, their activation times in the no-guidance test are comparable to the time in the adaptive-guidance test, suggesting that Group 2 participants have effectively mastered the gaze-depth interaction method without any visual guidance. Moreover, we also compare the activation time between the gaze-depth method (FocusFlow) and the baseline dwell method in Figure 18a. We could find that the two settings in FocusFlow achieve a similar activation time of around 1.3s, which is shorter than the 2-second dwell method, while the half-second dwell method achieved the shortest time.
Figure 18. Performance comparison between FocusFlow and dwell methods in different settings.
8.4.2. Failure Rate
We define ’failure rate’ as the ratio of successful selection attempts to the total number of attempts in each stage, considering an attempt a failure if it’s not completed within the time limit. Figure 18b reveals that all four settings exhibit a failure rate below 5%, which suggests that the reliability of FocusFlow is comparable to that of the traditional dwell method.
8.4.3. False Trigger Rate
We define the ’false trigger rate’ as the proportion of incorrect selections relative to the total number of selections, with ’incorrect selections’ being those where the wrong targets are chosen. As illustrated in Figure 18c, the half-second dwell method exhibits the highest false trigger rate. This may be attributed to the method’s overly sensitive nature, leading to an increased likelihood of unintentional target activations. In contrast, both settings in FocusFlow demonstrate a considerably lower false trigger rate, approximately 5%. This rate is marginally higher than that observed in the 2-second dwell method.
8.4.4. Overall Performance
In the preceding paragraphs, we evaluated various metrics of FocusFlow, comparing them against two baseline dwell time settings as depicted in Figure 18. Overall, FocusFlow demonstrates commendable performance with an average activation time of 1.3 seconds, a failure rate under 5%, and a relatively low false trigger rate of approximately 5%. While the half-second dwell method offers a shorter average activation time and a comparably low failure rate, its high incidence of false triggers could negatively impact user experience. Conversely, the 2-second dwell method, despite its stability and low rates of failure and false triggers, may lead to user frustration due to longer activation times, especially in certain scenarios. Additionally, the dwell method restricts users from continuously observing the target object, which might detract from the overall user experience. Thus, FocusFlow is a reliable and efficient interaction choice with an appropriate learning strategy applied to enhance the user experience in the VR space.
8.4.5. User Experience
Besides the quantitative metrics we discussed above, we also designed a Likert questionnaire to evaluate the user experience of FocusFlow, focusing on its usability, efficiency, comfort, overall preference, and potential applicability in various VR contexts, and compared the results with the dwell-based baseline method. The result in Figure 19 shows that FocusFlow receives positive feedback from participants, which proves its usability and learnability. Compared to the baseline method, FocusFlow is overall better and obviously outperforms it in terms of efficiency (Q.4, Q.5, Q.6) and overall experience (Q.9, Q.10). It also shows that the eye fatigue when using FocusFlow is more serious than the baseline dwell-based method, which should be an issue that needs to be emphasized and improved upon. In addition, users are wary of using hands-free gaze interaction as an alternative to other VR interaction methods. This suggests that gaze-only interaction cannot cope with all VR interaction scenarios, so compatibility and extendibility with other interaction methods should also be an important design consideration.
Figure 19. User Experience Likert Data.
9. Case Study
(a) Scene Overview
(b) Visual Cues
(c) Quick Preview Window
Figure 20. Case study in a shopping scenario.
FocusFlow facilitates intuitive and accurate hands-free selections in VR. To demonstrate how users interact with FocusFlow and their perceptions of the process, we present a case study set in a virtual fashion store shopping environment that mimics real-world scenarios. (Figure 19(a))
In this environment, each product can trigger a preview window that showcases its profile. This is consistent with the Virtual Window setting encountered during user training. Items within the store vary in distance from the user, ranging between 2 to 7 meters. Users are encouraged to freely explore these products, mirroring a genuine shopping experience. The dwell method and our depth-based method are employed successively to ascertain the experiential difference conferred by the addition of a depth input dimension.
Utilizing FocusFlow, users reported a seamless and efficient experience. Initially, they would pan their heads around to acquaint themselves with the virtual space. Gradually, their attention would focus on specific items of interest. At this juncture, visual cues become evident (Figure 19(b)). “The appearance of adaptive visual cues hinted at the target depth and gave me some feedback on my gaze depth, allowing me to confidently engage in depth adjustments to activate the virtual window.” “I didn’t experience significant visual disruption since the depth of the adaptive visual cues didn’t match the depth of the object I was observing. Thus, I could easily overlook its presence when focused on an item.” (P19) To access detailed information about an item, users adjusted their gaze to the depth of the virtual window (Figure 19(c)). This action promptly displayed the product details. Subsequently, by shifting their focus, users could effortlessly continue to survey the environment. “The activation and deactivation mechanism was very fluid. It felt as though a virtual window was consistently present; and whenever I looked at it, information would display.” (P8)
However, during free exploration using the dwell method, users faced considerable challenges. When their gaze lingered on a product, the sudden appearance of the information panel – triggered by surpassing the observation time threshold – obstructed their view. “The most significant flaw of this method is that I can’t freely gaze at an object indefinitely. I constantly have to be cautious to not stare too long, lest the window pops up, hindering my observation. I believe the freedom to observe the world undisturbed is crucial for user experience.” (P21) The lag between intention to inspect and the system’s response was also disruptive. “During that waiting period, even if it’s just a second, I felt anxious. It felt like wasted time since the system didn’t promptly capture my intent.” (P11) After inspecting a product, to exit the information panel, users had to move their gaze to its periphery, an action they deemed superfluous. “The dwell method required me to shift my gaze to the edge to exit. There was a continual need to oscillate between the center and periphery. In contrast, with the depth method, I merely had to naturally look past the translucent panel towards the distant environment to exit. Essentially, with the depth method, I wasn’t burdened with unnecessary actions since my goal was to resume observation, which in itself served as the exit mechanism.” (P9)
In summary, users welcomed the depth input dimension. “The depth method gave me a stronger sense of control. I could distinctly decide whether to activate a target.” (P6) “I preferred the depth method. Though time-based techniques are simpler, once I got the hang of the depth method, I noticed an improved experience.” (P8) However, they also highlighted that the depth method was more straining on the eyes after frequent use. “Compared to the dwell method, a downside of the depth approach is the fatigue it induces in my eyes due to the constant need for depth adjustments.” (P13) This feedback implies that the depth method might not be ideal for extended use in short durations, as continuous adjustments could tire the eyes.
10. Discussion and Future Work
10.1. User Fatigue and Cognitive Load
Our user studies demonstrate the effectiveness of visual cues combined with the learning process to master this new interaction input and eventually use the system without any visual cues. However, some users still appreciate the scaffolding that these cues provide and showed interest in using the cues as a continual depth perception guideline. Therefore, visual cues are not merely transitional aids and can be an integral aspect of the everyday usage. This is analogous to television subtitles, where viewers might prefer the use of subtitles despite their capability in comprehending the show without them as a way of reducing cognitive load and bolster comprehension.
10.2. Compatibility and Extendibility
The introduction of visual depth as a unique input dimension does not clash with other input modalities such as eye movements or blinks. Instead, FocusFlow can be seamlessly integrates with other gaze interactions. Moreover, the visual depth modality can be harmoniously coupled with other VR interaction modes like head and hand movements. For instance, a user might activate a virtual window using visual depth transition and subsequently interact with components within the window using hand gestures. Such rich and compound interactions are ripe avenues for exploration in future research.
10.3. Gaze Depth as a Continuous Input
The evaluations in Section 3.3 show that high-accuracy depth estimations are limited to close distances. While this limits the virtual windows to certain depths, we can expand this interaction method to a continuous input by tracking the continuous depth changes within a certain range. This will enable various new inputs such as scrolling on a webpage or rotating a 3D object by looking at micro-gaze depth changes. In addition, different mapping functions can be defined to match the gaze depth changes to the speed of interactions.
10.4. Virtual Window Depth Adaptability
To elevate user experience and reduce erroneous virtual window activation, the position of the interaction layer could be adapted to a target’s location. For example, when the target is a few meters away from the user, the interaction layer could be set to one meter in front of the user. However, if the target distance is less than one meter, the position of the interaction layer will be automatically set to half a meter to the user. However, such adaptability necessitates visual cues to inform users of the adjusted window position, ensuring seamless interactions. In addition, in this paper, we assumed that the virtual window depth is always displaces closer to the user than the interactive object. However, FocusFlow can be seamlessly extended to applications where virtual windows can be designed further from the interactive object. For example, in a surgical training application, the user can switch between observing different layers inside the patients body by actively changing their focal depth further from the upper surface of the body.
11. Conclusion
We present FocusFlow, a novel gaze-depth interaction technique that leverages the potential of visual depth as an intuitive input dimension. Our approach introduces a layer-based user interface that resonates with the concept of a “Virtual Window”, offering an efficient and hands-free selection method. By doing so, we address not only the traditional Midas Touch problem but also enhance the overall user experience in virtual reality (VR) interactions.
Considering that visual depth is a new input dimension for VR users, we place emphasis on the learnability of our design. To evaluate this aspect, we conducted both a user study and a case study to observe how novice users adapt to this new interaction method. The results demonstrate that gaze-depth interaction can be learned smoothly with adaptive visual cues that facilitate users’ perception of depth and develop muscle memory.
We hope this study will inspire more researchers to refine theoretical frameworks related to gaze-depth interaction. For instance, future work could focus on designing more effective learning processes, assisting users’ depth perception during transition periods, and integrating inputs from other modalities. These efforts would provide valuable theoretical support for promoting widespread adoption of this intuitive interaction method.
Appendix A Autostereogram
Figure 21 shows an autostereogram, a picture can show 3D patterns at a certain visual depth.
Figure 21. Autostereogram Example. When you put your gaze point sightly further and make the green circle and pink circle overlap, you can see a 3D image. Making the image as large as possible and not too close to the image will help you see the 3D image successfully.
Appendix B User Experience Questionnaire
1. Usability and Learnability
  • The gaze-based interaction method was easy to understand.
  • I quickly became comfortable using the gaze-based interaction method in VR.
  • The gaze-based interaction method was intuitive, requiring little or no explanation.
2. Responsiveness and Efficiency
  • The gaze-based interaction method was responsive to my eye movements.
  • The gaze-based interaction method allowed me to perform ”selection” operation efficiently.
  • I felt in control while using the gaze-based interaction method to perform ”selection” operation.
3. Comfort and Naturalness
  • The gaze-based interaction method felt natural and immersive.
  • I experienced minimal eye fatigue while using the gaze-based interaction method to perform ”selection” operation.
4. Preference and Overall Experience
  • I would prefer using the gaze-based interaction method over other VR interaction methods such as controllers.
  • The gaze-based interaction method enhanced my overall VR experience.
5. Applicability in Various Contexts
  • I believe the gaze-based interaction method could be useful in a variety of VR applications.